POTENTIAL ENERGY CURVES OF THE A^{1}Σ^{+} STATES OF AgH AND CuH

\underline{DOMINIQUE R.T. APPADOO}, ALIREZA SHAYESTEH, PETER F. BERNATH AND ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Some years ago Learner^{a} recorded electronic spectra involving the A^{1}Σ^{+} state of AgH, and found anomalous behaviour of the vibrational level spacings and B_{v} values which he attributed to the potential energy curve for this state having a “shelf”. More recently, Witek et al.^{b} reported ab initio results which supported Learner’s assertion that the potential energy function for this state had an unusual shape, but found that rather than have a shelf, there was an abrupt stiffening of the bond for vibrational levels above v ≈ 4 . In order to clarify this situation, we have combined new and previously recorded visible A−X spectra and high quality microwave and infrared data for ^{107/109}AgH and ^{107/109}AgD in a combined-isotopomer direct-potential-fit analysis which determines analytic potential energy and Born-Oppenheimer breakdown functions for this state. A similar approach is used to study the analogous A^{1}Σ^{+} state of ^{63/65}CuH and 63/65CuD.

a R.C.M. Learner, Proc. Roy. Soc. (London) A 269, 327 (1962).

b H.A. Witek, A. Viel and P.-O. Widmark, J. Chem. Phys. 116. 8396 (2002).